OpenStudy (anonymous):

he temperature decreased overnight by 15.2°C to a temperature of –12.8°C. Suppose t represents the temperature before it decreased. Which equation can be used to represent this situation? A. t – 15.2 = –12.8 B. t + 12.8 = –15.2 C. t – 12.8 = 15.2 D. t – (–15.2) = –12.8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after this i have one more but thanks for helping me

OpenStudy (anonymous):

No problem, ok, so it says that after it lost 15.2 Celsius it went down to -12.8; basically all we have to do is make an equation that says that T (temperature before decreasing) - 15.2 is equal to -12.8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[T-(-15.2) = -12.8\] A simplified version of the text I wrote up there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The answer is D, if you are still confused, tell me where and I'll help you out

OpenStudy (anonymous):

ok and i have one more

OpenStudy (anonymous):

Ok, best response me in this one, close this question and open another question whenever you are ready

OpenStudy (anonymous):

Dont forget to mention also

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok and On Wednesday, William wrote a check in the amount of $45.60. This brought his balance down to –$15.75. Suppose n represents William’s balance before he wrote the check. Which equation can be used to represent this situation? A. n + 15.75 = –45.60 B. n – 15.75 = 45.60 C. n + 45.60 = –15.75 D. n – 45.60 = –15.75